OpenStudy (anonymous):

find two numbers whose sum is 28 and whose difference is 20.

sammixboo (sammixboo):

24 + 4 = 28 24 - 4 = 20 :)

sammixboo (sammixboo):

I can't really explain how I got the answer, but that is what I got

sammixboo (sammixboo):

I think I just found the number in the middle of 20 and 28, and then subtracted that number from 28 to get 4. Take a look here. find two numbers whose sum is 100 and whose difference is 50 Number in the middle of 50 and 100 is 75 100 - 75 = 25 Now we have 75 + 25 = 100 75 - 25 = 50

sammixboo (sammixboo):

I don't know the proper way to solve this, ahaha

OpenStudy (anonymous):

yay chick

geerky42 (geerky42):

\(a+b = 28\) and \(a-b=20\) So add them together: \(a+b+a-b = 2a =48\\~\\a=24\) So \(a+b \to 24+b = 28\) \(b = 28-24 = 4\) \(\boxed{a=24,~b=4}\)

geerky42 (geerky42):

That's how you can get answer.

sammixboo (sammixboo):

Ah okay :)